# Characterization of Nanostructural Imperfections in Superconducting Quantum Circuits

### v1

@@ Acknowledgment
This work has received funding from the United Kingdom Engineering and Physical Sciences Research Council (EPSRC) under Grants No. EP/J013501/1, No. EP/N015118/1, and No. EP/T001062/1. M.B. acknowledges support from EPSRC QT Fellowship grant [EP/W027992/1]. S.C. acknowledges support from Schmidt Science.
We would like to acknowledge Superfab Nanofabrication Laboratory at Royal Holloway, University of London, and Optoelectronics Research Centre at University of Southampton where parts of device fabrication were performed as well as electron Physical Science Imaging Centre (ePSIC) at Diamond Light Source, where all materials characterizations were performed.

### v2

@@ Acknowledgment
This work has received funding from the United Kingdom Engineering and Physical Sciences Research Council (EPSRC) under Grants No. EP/N015118/1 and No. EP/T001062/1. M.B. acknowledges support from EPSRC QT Fellowship under Grant No. EP/W027992/1. S.C. acknowledges support from Schmidt Science.
We thank Diamond Light Source for access and support in use of the electron Physical Science Imaging Centre (Instrument E01 and E02 and proposal number MG37092) that contributed to the results presented here. We would like to acknowledge the Superfab Nanofabrication facility at Royal Holloway, University of London, and Optoelectronics Research Centre at University of Southampton where part of device fabrication was performed.